I Just dont get equasions i am just looking for an easier way to learn them than my teacher teaches me.
12-11a=45
18=-a+2
-8c+1=-3

Answer by Alan3354(69443)

I don't know what that teaching method might be.
----------
12-11a=45
The idea is to isolate the variable, ie, get it on one side of the equal sign by itself.
12-11a=45
Get rid of the 12 and the -11.
Subtract 12 (from both sides)
-11a = 33
Divide by the 11
-a = 3
Times -1
a = 3
Not a big deal.
-------------
You do these, any way you want. Just get the a (and the c) alone.
18=-a+2
-8c+1=-3